# Efficient Trajectory Forecasting and Generation with Conditional Flow Matching

### v1

@@ Abstract
Trajectory prediction and generation are vital for autonomous robots navigating dynamic environments. While prior research has typically focused on either prediction or generation, our approach unifies these tasks to provide a versatile framework and achieve state-of-the-art performance. Diffusion models, which are currently state-of-the-art for learned trajectory generation in long-horizon planning and offline reinforcement learning tasks, rely on a computationally intensive iterative sampling process. This slow process impedes the dynamic capabilities of robotic systems. In contrast, we introduce Trajectory Conditional Flow Matching (T-CFM), a novel data-driven approach that utilizes flow matching techniques to learn a solver time-varying vector field for efficient and fast trajectory generation. We demonstrate the effectiveness of T-CFM on three separate tasks: adversarial tracking, real-world aircraft trajectory forecasting, and long-horizon planning. Our model outperforms state-of-the-art baselines with an increase of 35% in predictive accuracy and 142% increase in planning performance. Notably, T-CFM achieves up to 100$\times$ speed-up compared to diffusion-based models without sacrificing accuracy, which is crucial for real-time decision making in robotics.
@@ II-B Learning-Based Approaches
Recent works in trajectory prediction for various domains, such as aircraft navigation (FlightBERT) [10], social navigation [3], and autonomous driving [18], have employed supervised learning methods. These approaches often utilize autoregressive models and log-likelihood based training to learn predictive models from data. Graph-based Adversarial Modeling with Mutual Information (GrAMMI) [26] is a recent framework that explicitly models a multimodal distribution using a combination of a Gaussian Mixture Model regularized by mutual information. While effective in certain scenarios, these models can struggle with capturing long-horizon multi-modal distributions, which are common in real-world trajectory data. We leverage a generative modeling approach rather than a supervised learning one, enabling more diverse multimodal outputs for accurate prediction in complex real-world scenarios.
@@
Generative modeling techniques, which have shown great success in computer vision tasks [20], provide a promising avenue to augment trajectory prediction by learning to model complex, multimodal distributions. Recently, diffusion-based probabilistic models [20] have dominated many generative modeling tasks. Diffusion models generate samples by iteratively denoising a Gaussian distribution, allowing them to capture complex, multimodal distributions. Diffuser [11], a diffusion based approach for learning trajectories, is state of the art in long horizon planning and offline reinforcement learning. Constrained Agent-based Diffusion for Enhanced Multi-Target Tracking (CADENCE) [25], similarly extends the Diffuser framework for Adversarial Tracking and modeling multi-agent behaviors. The main drawback of diffusion models is their iterative denoising process, which can be computationally slow, limiting their real-time applicability.
@@ IV-A Flow Matching Formulation
We assume a Gaussian flow between $\tau_{0}$ and $\tau_{1}$ with standard deviation $\sigma$ and model the probability path $p_{t}(\tau|z)$ and vector field $u_{t}(\tau|z)$ as shown in Equations IV-A. Equation 3a defines the probability path as a Gaussian distribution centered at a linear interpolation between $\tau_{0}$ and $\tau_{1}$ at time $t$. In the top portion of Figure 2, we show a visualization of the linear interpolation used to generate intermediate trajectories between $\tau_{0}$ and $\tau_{1}$. Equation 3b defines the target vector field simply as the difference vector pointing from the starting point $\tau_{0}$ to the end point $\tau_{1}$.
@@
4:     $t\sim U(0,1)$triangleright Draw timestep
5:     $\tau_{1},c\sim\mathcal{D}$triangleright Draw target trajectory and conditioning
6:     $\tau\sim p_{t}(\tau|z)=\mathcal{N}\left(\tau|t\tau_{1}+(1-t)\tau_{0},\sigma^{2}\right)$triangleright Eq. 3a
7:     $u_{t}(\tau|z)=\tau_{1}-\tau_{0}$triangleright Eq. 3b
8:     $\mathcal{L}_{CFM}(\theta)=||v_{\theta}(t,\tau,c)-u_{t}(\tau|z)||^{2}$triangleright Loss
9:     $\theta=\theta+\alpha\nabla_{\theta}\mathcal{L}_{CFM}(\theta)$triangleright Update Model
@@
The neural network model used to parameterize $v_{\theta}(t,\tau,c)$ is a 1D Convolutional Temporal U-Net based on prior diffusion work [11, 25]. 1D convolutions slide over the time dimension of the input trajectory $\tau$, capturing temporal patterns and dependencies without being autoregressive. This allows for efficient parallel processing of the entire trajectory. The model also incorporates Feature-Wise Linear Modulation (FiLM) Layers [15] to condition the model with relevant context information, $c$. For each domain, we will describe the context vector ($c$) used in Section V. By using the same base architecture as prior diffusion work, we demonstrate that our training methodology generates better models irrespective of model parameter count and architecture.
@@ IV-B Sampling
Given a trained flow model $v_{\theta}(t,\tau,c)$, the sampling procedure utilizes an ODE solver to recover the solution to Equation 1. We can denote the solution of the ODE with $\phi_{t}(\tau)$, where $\phi_{0}(\tau)=\tau$ and $\phi_{t}(\tau)$ is the transformation of our trajectory $\tau$ transported along the vector field from time 00 to time $t$. In Algorithm 2, we show the sampling procedure using the Euler method (Line 9) but any off the shelf ODE solver can be used. In our experiments, we choose to use the Euler sampling method as the number of sampling steps is easily adjustable. The bottom portion of Figure 2 shows how the sampling procedure moves from a prior noisy trajectory $\tau_{0}$ to a trajectory that lies within the data distribution $\tau_{1}$.
@@
1:Samplable $q(\tau_{0})=\mathcal{N}(0,I)$, trained flow network $v_{\theta}(t,\tau,c)$, number of sampling timesteps $N$
@@
6:         $\tau_{t/N}^{0}\leftarrow s^{0}$triangleright Set Start State
7:         $\tau_{t/N}^{h}\leftarrow s^{h}$triangleright Set Goal State
@@
The key feature of parameterizing the probability flow and vector field through Equation IV-A is that it enables our model to learn straight flows as compared to diffusion models. In diffusion models, the sampling process involves gradually denoising a Gaussian noise sample over many steps, following a complex path in the data space. This typically requires a large number of sampling steps to generate high-quality samples. In contrast, our flow matching approach learns a direct, straight path from the starting noise distribution to the target data distribution by modeling the probability path as a linear interpolation between the starting and ending points (Equation IV-A), encouraging the model to find the most efficient trajectory that matches the true data distribution.
@@ V Evaluation and Domains
We test our model in three different tasks and domains: 1) Adversarial Tracking 2) Trajectory Forecasting, and 3) Long-Horizon Planning. These domains test our model’s capability of generating accurate multimodal trajectory predictions and plans for robots to use. Visualizations for the training data and domains are shown in Figure 3.
@@ V-A Adversarial Tracking
Adversarial tracking aims to predict an adversary’s future trajectory $\tau$ given past historical information, $c$. These domains are challenging due to the adversary’s potential multiple strategies and the observers’ often incomplete or sparse data. We assess our flow-matching tracking models using the Prison Escape and Smuggler scenarios, as introduced in previous work [26].
@@ VI-A1 Tracking Capabilities of Flow Matching
P-High VRNN 0.105 0.059 0.100 0.117 0.145
@@
NTI-Low VRNN 0.147 0.156 0.186 0.187 0.203
GrAMMI 0.121 0.144 0.181 0.183 0.193
CADENCE 0.112 0.123 0.135 0.148 0.160
T-CFM (Ours) 0.110 0.121 0.133 0.145 0.156
NTI-High VRNN 0.138 0.153 0.183 0.179 0.185
GrAMMI 0.131 0.163 0.174 0.175 0.184
CADENCE 0.113 0.124 0.138 0.152 0.163
T-CFM (Ours) 0.112 0.125 0.137 0.150 0.162
TABLE I: Average Displacement Error Results for three Prisoner Escape (P-low, P-med, P-high) and two Narco Traffic Interdiction (NTI-low, NTI-high) Datasets. Bolded values represent the best performing model.
We report our results on the three Prisoner Escape datasets and two Narco Interdiction datasets in Table I. We show that T-CFM outperform or matches the prior baselines on all prediction horizons with the greatest advantages on the Prisoner-Medium and Prisoner-High datasets, showcasing a 17% and 12% increase in predictive accuracy respectively. We hypothesize that the flow-matching models are able to better incorporate the dense detection history information than the diffusion models because the flow field is deterministic and does not include adding an additional noise component. This may benefit the flow matching models to generate more confident and correct trajectories as compared to the diffusion models.
@@ 00 $15$ $30$ 00 $15$ $30$ 00 $15$ $30$
TABLE II: Aircraft Trajectory Forecasting: T-CFM achieves lower MAE and RMSE on Latitude, Longitude, and Altitude.
@@ VI-B Aircraft Trajectory Forecasting
Table II shows our model’s performance as compared to FlightBERT. We find that our method outperforms FlightBERT on all metrics with an average improvement of $35.4\%$ over all metrics. We hypothesize two main reasons that our method outperforms FlightBERT. First, our method is not autoregressive generates the whole trajectory at once. This provides an advantage as errors may not accumulate over the prediction time horizon. FlightBERT was only tested for shorter horizon predictions. Meanwhile, we are interested in longer horizon predictions of up to 30 minutes compared to the 5 minute horizon for the dataset in FlightBERT. Second, we hypothesize that the multimodal capabilities of our model is important for our flight trajectory dataset. Unlike commercial flights, the behavior of the Cessna aircraft does not travel in straight paths and consists of multiple heading changes throughout its path. We show that T-CFM better models these diverse trajectories than prior work.
@@ VI-C Long-Horizon Planning through Imitation Learning
We compare how well our models perform against Diffuser [11], the current state-of-the-art method for learning how to plan solely from data on the Maze2D task (Table III). We find that with just a single sampling step, our method significantly outperforms the diffusion models, achieving a 142% increase in score. Similar to our analysis in Adversarial Tracking, the linear probability flows allows us to immediately infer high quality samples. Additionally, we provide a visualization of the sampling procedure with just two sampling timesteps in Figure 5. Here we show that Diffuser produces a plan that intersects with the wall, as it is requires a large number of sampling steps to produce coherent plans. Meanwhile our method plans a collision free path. Additionally, the middle trajectory $\tau_{1/2}$ shows more coherence and has less spread in T-CFM than Diffuser. This supports the hypothesis that our T-CFM’s flow field is more efficient at transforming noisy trajectories into realistic ones.
@@
Maze2D U-Maze 50.7$\pm$6.7 106.7$\pm$2.7 112.5$\pm$11.2 122.1$\pm$1.4
Maze2D Medium 21.7$\pm$13.5 112.2$\pm$1.5 123.3$\pm$1.6 123.8$\pm$3.5
Maze2D Large 30.3$\pm$6.9 111.0$\pm$13.5 112.6$\pm$16.3 104.3$\pm$3.4
@@
Multi2D U-Maze 69.8$\pm$16.0 129.8$\pm$3.0 127.3$\pm$3.3 129.5$\pm$0.9
Multi2D Medium 58.4$\pm$5.3 116.5$\pm$2.8 124.2$\pm$1.2 126.5$\pm$4.1
Multi2D Large 35.8$\pm$4.2 121.7$\pm$5.0 138.7$\pm$5.9 127.3$\pm$8.6
@@
TABLE III: The performance of T-CFM and Diffuser on the long-horizon Maze2D compared when given a single sampling timestep $N=1$ and maximum sampling timesteps, $N=256$. Our model (T-CFM) is able to drastically reduce the number of sampling steps required to generate feasible plans whereas the Diffuser model fails at $N=1$.
@@ VII Limitations and Future Work
We currently do not explicitly reason about potential interaction between multiple agents. Future work includes extending our framework to other domains such as social navigation and trajectory prediction in autonomous driving, which require explicit reasoning of other agents. Additionally, further work could be used to increase the expressiveness of our flow matching models. Recent work on stochastic bridge matching [23] could be used to combine the deterministic ODE formulation of flow matching with the stochastic SDE formulation in diffusion models. Finally, future experiments on using T-CFM in dynamic situations can better show our model’s capabilities for real robot tasks.

### v2

@@ Abstract
[…]
@@ II-B Learning-Based Approaches
Recent works in trajectory prediction for various domains, such as aircraft navigation (FlightBERT) [11], social navigation [4], and autonomous driving [19], have employed supervised learning methods. These approaches often utilize autoregressive models and log-likelihood based training to learn predictive models from data. Graph-based Adversarial Modeling with Mutual Information (GrAMMI) [27] is a recent framework that explicitly models a multimodal distribution using a combination of a Gaussian Mixture Model regularized by mutual information. While effective in certain scenarios, these models can struggle with capturing long-horizon multi-modal distributions, which are common in real-world trajectory data. We leverage a generative modeling approach rather than a discriminative one, enabling more diverse multimodal outputs for accurate prediction in complex real-world scenarios.
@@
Generative modeling techniques, which have shown great success in computer vision tasks [21], provide a promising avenue to augment trajectory prediction by learning to model complex, multimodal distributions. Recently, diffusion-based probabilistic models [21] have dominated many generative modeling tasks. Diffusion models generate samples by iteratively denoising a Gaussian distribution, allowing them to capture complex, multimodal distributions. Diffuser [12] and Motion Planning Diffusion [3], are diffusion based approaches for learning trajectories, representing the state of the art in long horizon planning and offline reinforcement learning. Constrained Agent-based Diffusion for Enhanced Multi-Target Tracking (CADENCE) [26], similarly extends the Diffuser framework for Adversarial Tracking and modeling multi-agent behaviors. The main drawback of diffusion models is their iterative denoising process, which can be computationally slow, limiting their real-time applicability.
@@ IV-A Flow Matching Formulation
We assume a Gaussian flow between $\tau_{0}$ and $\tau_{1}$ with standard deviation $\sigma$ and model the probability path $p_{t}(\tau|z)$ and vector field $u_{t}(\tau|z)$ as shown in Equations 3. Equation 3a defines the probability path as a Gaussian distribution centered at a linear interpolation between $\tau_{0}$ and $\tau_{1}$ at time $t$. In the top portion of Figure 2, we show a visualization of the linear interpolation used to generate intermediate trajectories between $\tau_{0}$ and $\tau_{1}$. Equation 3b defines the target vector field simply as the difference vector pointing from the starting point $\tau_{0}$ to the end point $\tau_{1}$.
@@
4:  $t\sim U(0,1)$ $\triangleright$ Draw timestep
5:  $\tau_{1},c\sim\mathcal{D}$ $\triangleright$ Draw target trajectory and conditioning
6:  $\tau\sim p_{t}(\tau|z)=\mathcal{N}\left(\tau|t\tau_{1}+(1-t)\tau_{0},\sigma^{2}\right)$ $\triangleright$ Eq. 3a
7:  $u_{t}(\tau|z)=\tau_{1}-\tau_{0}$ $\triangleright$ Eq. 3b
8:  $\mathcal{L}_{CFM}(\theta)=||v_{\theta}(t,\tau,c)-u_{t}(\tau|z)||^{2}$ $\triangleright$ Loss
9:  $\theta=\theta+\alpha\nabla_{\theta}\mathcal{L}_{CFM}(\theta)$ $\triangleright$ Update Model
@@
The neural network model used to parameterize $v_{\theta}(t,\tau,c)$ is a 1D Convolutional Temporal U-Net based on prior diffusion work [12, 26]. 1D convolutions slide over the time dimension of the input trajectory $\tau$, capturing temporal patterns and dependencies without being autoregressive. This allows for efficient parallel processing of the entire trajectory. The model also incorporates Feature-Wise Linear Modulation (FiLM) Layers [16] to condition the model with relevant context information, $c$. For each domain, we will describe the context vector ($c$) used (Section V). By using the same base architecture as prior diffusion work, we demonstrate that our training methodology generates better models irrespective of model parameter count and architecture.
@@ IV-B Sampling
Given a trained flow model $v_{\theta}(t,\tau,c)$, the sampling procedure utilizes an ODE solver to recover the solution to Equation 1. We can denote the solution of the ODE with $\phi_{t}(\tau)$, where $\phi_{0}(\tau)=\tau$ and $\phi_{t}(\tau)$ is the transformation of our trajectory $\tau$ transported along the vector field from time 0 to time $t$. In Algorithm 2, we show the sampling procedure using the Euler method (Line 9) but any off the shelf ODE solver can be used. In our experiments, we choose to use the Euler sampling method as the number of sampling steps is easily adjustable. The bottom portion of Figure 2 shows how the sampling procedure moves from a prior noisy trajectory $\tau_{0}$ to a trajectory that lies within the data distribution $\tau_{1}$.
@@
1:Samplable $q(\tau_{0})=\mathcal{N}(0,I)$, trained flow network $v_{\theta}(t,\tau,c)$, number of sampling timesteps NN
@@
6:   $\tau_{t/N}^{0}\leftarrow s^{0}$ $\triangleright$ Set Start State
7:   $\tau_{t/N}^{h}\leftarrow s^{h}$ $\triangleright$ Set Goal State
@@
The key feature of parameterizing the probability flow and vector field through Equation 3 is that it enables our model to learn straight flows as compared to diffusion models. In diffusion models, the sampling process involves gradually denoising a Gaussian noise sample over many steps, following a complex path in the data space. This typically requires a large number of sampling steps to generate high-quality samples. In contrast, our flow matching approach learns a direct, straight path from the starting noise distribution to the target data distribution by modeling the probability path as a linear interpolation between the starting and ending points (Equation 3), encouraging the model to find the most efficient trajectory that matches the true data distribution.
@@ V Evaluation and Domains
We test our model in three different tasks and domains: 1) Adversarial Tracking 2) Trajectory Forecasting, and 3) Long-Horizon Planning. These domains test our model’s capability of generating accurate multimodal trajectory predictions and plans for robots to use. Visualizations for the training data and domains are shown in Figure 3 and a summary of the states and context vectors used are in Table I.
Domain State in Trajectory ($\tau$) Context Vector ($c$)
Adversarial Tracking Position (x, y) Historical detection
coordinates information
Aircraft Trajectory Longitude, Latitude, 5-minute history
Altitude of past states
Long-Horizon Planning Position (x, y) Start and goal
coordinates states
TABLE I: Summary of State and Context Vectors for Different Domains
@@ V-A Adversarial Tracking
Adversarial tracking aims to predict an adversary’s future trajectory $\tau$ given past historical information, $c$. These domains are challenging due to the adversary’s potential multiple strategies and the observers’ often incomplete or sparse data. We assess our flow-matching tracking models using the Prison Escape scenarios, as introduced in previous work [27].
@@ VI-A1 Tracking Capabilities of Flow Matching
P-High Particle Filter 0.041 0.084 0.102 0.119 0.133
VRNN 0.105 0.059 0.100 0.117 0.145
@@
TABLE II: Average Displacement Error Results for three Prisoner Escape (P-low, P-med, P-high) Datasets. Bolded values represent the best performing model.
We report our results on the three Prisoner Escape datasets in Table II. We show that T-CFM outperform or matches the prior baselines on all prediction horizons with the greatest advantages on the Prisoner-Medium and Prisoner-High datasets, showcasing a 17% and 12% increase in predictive accuracy respectively. We hypothesize that the flow-matching models are able to better incorporate the dense detection history information than the diffusion models because the flow field is deterministic and does not include adding an additional noise component. This may benefit the flow matching models to generate more confident and correct trajectories as compared to the diffusion models.
@@ VI-A2 Sampling Speed Analysis
TABLE III: Aircraft Trajectory Forecasting: T-CFM achieves lower MAE and RMSE on Latitude, Longitude, and Altitude.
@@ VI-B Aircraft Trajectory Forecasting
Table III shows our model’s performance as compared to FlightBERT. We find that our method outperforms FlightBERT on all metrics with an average improvement of $35.4\%$ over all metrics. We hypothesize two main reasons that our method outperforms FlightBERT. First, our method is not autoregressive generates the whole trajectory at once. This provides an advantage as errors may not accumulate over the prediction time horizon. FlightBERT was only tested for shorter horizon predictions. Meanwhile, we are interested in longer horizon predictions of up to 30 minutes compared to the 5 minute horizon for the dataset in FlightBERT. Second, we hypothesize that the multimodal capabilities of our model is important for our flight trajectory dataset. Unlike commercial flights, the behavior of the Cessna aircraft does not travel in straight paths and consists of multiple heading changes throughout its path. We show that T-CFM better models these diverse trajectories than prior work.
@@ VI-C Long-Horizon Planning through Imitation Learning
We compare how well our models perform against Diffuser [12], the current state-of-the-art method for learning how to plan solely from data on the Maze2D task (Table IV). We find that with just a single sampling step, our method significantly outperforms the diffusion models, achieving a 142% increase in score. Similar to our analysis in Adversarial Tracking, the linear probability flows allows us to immediately infer high quality samples. Additionally, we provide a visualization of the sampling procedure with just two sampling timesteps in Figure 5. Here we show that Diffuser produces a plan that intersects with the wall, as it is requires a large number of sampling steps to produce coherent plans. Meanwhile our method plans a collision free path. Additionally, the middle trajectory $\tau_{1/2}$ shows more coherence and has less spread in T-CFM than Diffuser. This supports the hypothesis that our T-CFM’s flow field is more efficient at transforming noisy trajectories into realistic ones.
@@
[…]
@@
[…]
@@
TABLE IV: The performance of T-CFM and Diffuser on the long-horizon Maze2D compared when given a single sampling timestep $N=1$ and maximum sampling timesteps, $N=256$. Our model (T-CFM) is able to drastically reduce the number of sampling steps required to generate feasible plans whereas the Diffuser model fails at $N=1$.
@@ VII Limitations and Future Work
Our current approach does not explicitly consider multi-agent interactions. Future work includes extending to social navigation and autonomous driving scenarios, which require reasoning about other agents. We aim to increase our flow matching models’ expressiveness, potentially incorporating stochastic bridge matching [24] to combine deterministic ODE and stochastic SDE formulations. Further experiments with T-CFM in dynamic situations will better demonstrate its capabilities for real-world robotic tasks.